# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ 4.7 Summarization
The evaluation results are presented in Table 13, we observe that our model, EMMA-500, performs comparably to other Llama2-based models like TowerInstruct Llama 2 7B. While EMMA-500 shows strength in certain language categories, it does not significantly outperform these models overall. This suggests that, although EMMA-500 is effective in multilingual summarization tasks, there is room for improvement to achieve more consistent performance across all languages. It is also important to note that this test set includes only 44 languages, limiting the evaluation of EMMA-500’s capabilities in low-resource languages.
Table 13: Results on XL-Sum (ROUGE-L/BERTScore).
@@
Llama 2 7B 0.07/0.67 0.07/0.67 0.07/0.66 0.06/0.65 0.10/0.62 0.08/0.71
Llama 2 7B Chat 0.09/0.68 0.09/0.68 0.08/0.68 0.08/0.67 0.12/0.64 0.10/0.74
CodeLlama 2 7B 0.07/0.66 0.07/0.65 0.07/0.66 0.06/0.63 0.10/0.64 0.08/0.71
LLaMAX Llama 2 7B 0.05/0.65 0.05/0.65 0.05/0.65 0.05/0.61 0.07/0.62 0.06/0.69
LLaMAX Llama 2 7B Alpaca 0.10/0.69 0.10/0.70 0.09/0.69 0.09/0.68 0.13/0.65 0.12/0.74
MaLA-500 Llama 2 10B v1 0.05/0.64 0.06/0.64 0.05/0.64 0.05/0.61 0.07/0.64 0.06/0.68
MaLA-500 Llama 2 10B v2 0.05/0.64 0.06/0.64 0.05/0.64 0.05/0.61 0.07/0.64 0.06/0.69
Yayi Llama 2 7B 0.08/0.67 0.09/0.67 0.07/0.67 0.07/0.66 0.11/0.61 0.09/0.71
TowerBase Llama 2 7B 0.08/0.67 0.08/0.67 0.07/0.67 0.07/0.65 0.10/0.63 0.09/0.71
TowerInstruct Llama 2 7B 0.09/0.68 0.09/0.69 0.09/0.69 0.08/0.66 0.12/0.64 0.09/0.73
Occiglot Mistral 7B v0.1 0.07/0.66 0.08/0.66 0.07/0.66 0.07/0.64 0.10/0.63 0.09/0.71
Occiglot Mistral 7B v0.1 Instruct 0.08/0.67 0.09/0.67 0.08/0.67 0.08/0.64 0.10/0.64 0.09/0.73
BLOOM 7B 0.07/0.65 0.07/0.65 0.07/0.65 0.06/0.62 0.10/0.58 0.08/0.71
BLOOMZ 7B 0.11/0.70 0.14/0.69 0.09/0.70 0.10/0.69 0.12/0.66 0.15/0.74
mGPT 0.04/0.60 0.05/0.60 0.04/0.60 0.02/0.56 0.07/0.59 0.04/0.63
mGPT-13B 0.05/0.62 0.06/0.62 0.05/0.63 0.04/0.59 0.08/0.59 0.05/0.67
Yayi 7B 0.12/0.70 0.14/0.69 0.10/0.69 0.11/0.69 0.12/0.66 0.15/0.75
Llama 3 8B 0.08/0.67 0.08/0.67 0.08/0.67 0.08/0.65 0.10/0.65 0.11/0.72
Llama 3.1 8B 0.09/0.67 0.08/0.66 0.08/0.67 0.08/0.65 0.09/0.66 0.10/0.72
Gemma 2 9B 0.07/0.65 0.07/0.64 0.07/0.66 0.07/0.63 0.09/0.64 0.09/0.71
Gemma 7B 0.07/0.65 0.07/0.63 0.06/0.65 0.07/0.67 0.07/0.60 0.08/0.63
Qwen 1.5 7B 0.10/0.69 0.11/0.69 0.09/0.69 0.10/0.68 0.11/0.65 0.11/0.74
Qwen 2 7B 0.10/0.69 0.12/0.69 0.09/0.69 0.10/0.68 0.11/0.65 0.11/0.74
EMMA-500 Llama 2 7B 0.09/0.67 0.08/0.66 0.08/0.67 0.08/0.69 0.11/0.66 0.10/0.67
4.8 Natural Language Inference
Table 14: 0-shot results on XNLI (ACC).
Model Avg High Medium Low
Llama 2 7B 0.4019 0.4526 0.3772 0.3497
Llama 2 7B Chat 0.3858 0.4277 0.3675 0.3387
CodeLlama 2 7B 0.4019 0.4627 0.3729 0.3386
LLaMAX Llama 2 7B 0.4427 0.4653 0.4264 0.4303
LLaMAX Llama 2 7B Alpaca 0.4509 0.4847 0.4280 0.4289
MaLA-500 Llama 2 10B v1 0.3811 0.4210 0.3585 0.3465
MaLA-500 Llama 2 10B v2 0.3811 0.4210 0.3585 0.3465
YaYi Llama 2 7B 0.4128 0.4732 0.3841 0.3494
TowerBase Llama 2 7B 0.3984 0.4608 0.3633 0.3439
TowerInstruct Llama 2 7B 0.4036 0.4707 0.3692 0.3379
Occiglot Mistral 7B v0.1 0.4235 0.4990 0.3839 0.3519
Occiglot Mistral 7B v0.1 Instruct 0.4081 0.4758 0.3718 0.3452
BLOOM 7B 0.4160 0.4513 0.3969 0.3838
BLOOMZ 7B 0.3713 0.4002 0.3556 0.3451
mGPT 0.4051 0.4297 0.3965 0.3730
XGLM 7.5B 0.4375 0.4572 0.4216 0.4300
YaYi 7B 0.3987 0.4385 0.3824 0.3515
Llama 3 8B 0.4497 0.4882 0.4384 0.3956
Llama 3.1 8B 0.4562 0.4961 0.4404 0.4083
Gemma 7B 0.4258 0.4644 0.4100 0.3801
Gemma 2 9B 0.4674 0.4850 0.4511 0.4649
Qwen 2 7B 0.4277 0.4731 0.4135 0.3653
Qwen 1.5 7B 0.3947 0.4095 0.3880 0.3783
EMMA-500 Llama 2 7B 0.4514 0.4609 0.4440 0.4471
@@ 4.9 Math
We evaluate the ability of LLMs to solve grade-school math problems across multiple languages on the MGSM (Multilingual Grade School Math) benchmark (112). MGSM is an extension of the GSM8K (Grade School Math 8K) dataset (25) by translating 250 of the original GSM8K problems into ten languages. We also split these ten languages into three groups, i.e., high-resource (Spanish, French, German, Russian, Chinese, and Japanese), medium-resource (Thai, Swahili, and Bengali), and low-resource (Telugu). Table 15 shows the results for 3-shot prompting with a flexible match to obtain the answers in model generation. We evaluate all the models by directly prompting the questions (denoted as direct) and the questions with answers followed by Chain-of-Thoughts prompt in the same languages as the subset being evaluated (denoted as CoT) (112). The results show that Llama 2 7B is a weak model in the MGSM math task. The base model and its variants failed in most settings. Multilingual LLMs such as BLOOM, XGLM, and YaYi are also subpar at this task. Recent LLMs like Llama 3, Qwen, and Gemma obtain reasonable performance, and Qwen series models are the best. Our model improves the Llama 2 7B model remarkably in both prompt strategies. For direct prompting, our model has an average accuracy of 0.1702, 7% higher than the Llama 2 7B chat model. For CoT-based prompting, our model increases the score of the Llama 2 7B chat model from 0.1 to 0.3 (20% higher) and slightly outperforms Llama 3 and 3.1 8B models.
Table 15: 3-shot results (ACC) on MGSM obtained by direct and CoT prompting.
@@
Llama 2 7B 0.0669 0.0807 0.0213 0.0120 0.0636 0.0760 0.0213 0.0080
Llama 2 7B Chat 0.1022 0.1373 0.0213 0.0080 0.1091 0.1353 0.0280 0.0160
CodeLlama 2 7B 0.0593 0.0707 0.0293 0.0120 0.0664 0.0873 0.0267 0.0200
LLaMAX Llama 2 7B 0.0335 0.0400 0.0200 0.0080 0.0362 0.0433 0.0227 0.0240
LLaMAX Llama 2 7B Alpaca 0.0505 0.0520 0.0400 0.0160 0.0635 0.0807 0.0413 0.0080
MaLA-500 Llama 2 10B v1 0.0091 0.0133 0.0027 0.0000 0.0073 0.0127 0.0000 0.0000
MaLA-500 Llama 2 10B v2 0.0091 0.0133 0.0027 0.0000 0.0073 0.0127 0.0000 0.0000
TowerBase Llama 2 7B 0.0615 0.0833 0.0173 0.0080 0.0616 0.0860 0.0240 0.0080
TowerInstruct Llama 2 7B 0.0724 0.0953 0.0173 0.0200 0.0824 0.1047 0.0187 0.0120
Occiglot Mistral 7B v0.1 0.1331 0.1687 0.0453 0.0160 0.1407 0.1880 0.0360 0.0120
Occiglot Mistral 7B v0.1 Instruct 0.2276 0.2980 0.0747 0.0280 0.2216 0.3040 0.0787 0.0280
BLOOM 7B 0.0287 0.0260 0.0280 0.0360 0.0229 0.0220 0.0147 0.0200
BLOOMZ 7B 0.0255 0.0267 0.0240 0.0120 0.0215 0.0167 0.0307 0.0200
mGPT 0.0135 0.0167 0.0053 0.0000 0.0142 0.0193 0.0067 0.0000
mGPT 13B 0.0131 0.0180 0.0067 0.0000 0.0153 0.0167 0.0120 0.0000
XGLM 7.5B 0.0102 0.0120 0.0067 0.0200 0.0116 0.0107 0.0120 0.0280
Yayi 7B 0.0276 0.0293 0.0147 0.0240 0.0302 0.0293 0.0240 0.0200
Llama 3 8B 0.2745 0.2787 0.2613 0.0560 0.2813 0.2853 0.2667 0.0520
Llama 3.1 8B 0.2836 0.2900 0.2613 0.0440 0.2731 0.2727 0.2547 0.0840
Gemma 7B 0.3822 0.3660 0.3827 0.2720 0.3578 0.3467 0.3707 0.2680
Gemma 2 9B 0.3295 0.2800 0.3573 0.3080 0.4469 0.3607 0.5200 0.4720
Qwen 2 7B 0.4895 0.5440 0.3880 0.1480 0.4469 0.3607 0.5200 0.4720
Qwen 1.5 7B 0.3156 0.4000 0.1600 0.0400 0.5147 0.5893 0.3907 0.1440
EMMA-500 Llama 2 7B 0.1702 0.1920 0.1187 0.0240 0.3036 0.4060 0.1480 0.0240
@@ 4.10 Machine Reading Comprehension
BELEBELE (14) is a machine reading comprehension dataset covering 122 languages including high- and low-resource languages. Each question offers four multiple-choice answers based on a short passage from the FLORES-200 dataset. This benchmark is very challenging, even the English version of it presents remarkable challenges for advanced models. Table 16 shows the zero-shot results in different resource groups. Our continual pre-training improves the Llama 2 7B base model. But Llama 2 7 B-based models mostly fail in this task and get quasi-random results. Recent advanced models like Llama 3.1 and Qwen 2 get reasonable results.
Table 16: 0-shot results (ACC) on BELEBELE.
@@
Llama 2 7B 0.2627 0.2676 0.2635 0.2607 0.2641 0.2627
Llama 2 7B Chat 0.2905 0.3184 0.2997 0.2895 0.2947 0.2909
CodeLlama 2 7B 0.2738 0.2737 0.2733 0.2730 0.2730 0.2738
LLaMAX Llama 2 7B 0.2309 0.2323 0.2315 0.2307 0.2310 0.2308
LLaMAX Llama 2 7B Alpaca 0.2448 0.2546 0.2482 0.2441 0.2460 0.2449
MaLA-500 Llama 2 10B v1 0.2296 0.2302 0.2298 0.2297 0.2298 0.2297
MaLA-500 Llama 2 10B v2 0.2296 0.2302 0.2298 0.2297 0.2298 0.2297
YaYi Llama 2 7B 0.2832 0.2964 0.2867 0.2811 0.2837 0.2826
TowerBase Llama 2 7B 0.2636 0.2743 0.2685 0.2629 0.2648 0.2634
TowerInstruct Llama 2 7B 0.2793 0.2988 0.2851 0.2757 0.2819 0.2792
Occiglot Mistral 7B v0.1 0.3016 0.3225 0.3094 0.3002 0.3040 0.3015
Occiglot Mistral 7B v0.1 Instruct 0.3205 0.3414 0.3262 0.3174 0.3240 0.3208
BLOOM 7B 0.2411 0.2425 0.2452 0.2412 0.2411 0.2408
BLOOMZ 7B 0.3932 0.4543 0.4367 0.4151 0.4008 0.3951
mGPT 0.2396 0.2401 0.2381 0.2400 0.2394 0.2398
XGLM 7.5B 0.2485 0.2552 0.2517 0.2489 0.2490 0.2485
YaYi 7B 0.3797 0.4437 0.4271 0.4049 0.3872 0.3809
Llama 3 8B 0.4073 0.4607 0.4292 0.4103 0.4187 0.4088
Llama 3.1 8B 0.4519 0.5250 0.4801 0.4565 0.4674 0.4534
Gemma 7B 0.4337 0.5263 0.4783 0.4482 0.4543 0.4394
Gemma 2 9B 0.5449 0.6410 0.5890 0.5583 0.5685 0.5505
Qwen 2 7B 0.4931 0.5762 0.5220 0.5004 0.5116 0.4948
Qwen 1.5 7B 0.4183 0.4886 0.4479 0.4218 0.4300 0.4178
EMMA-500 Llama 2 7B 0.2675 0.2832 0.2818 0.2758 0.2714 0.2694
@@
Table 17: 5-shot results (ACC) on ARC multilingual.
@@
Llama 2 7B 0.2756 0.3312 0.2731 0.2102
Llama 2 7B Chat 0.2802 0.3369 0.2779 0.2129
CodeLlama 2 7B 0.2523 0.2886 0.2464 0.2165
LLaMAX Llama 2 7B 0.2609 0.3000 0.2592 0.2148
LLaMAX Llama 2 7B Alpaca 0.3106 0.3689 0.3185 0.2249
MaLA-500 Llama 2 10B v1 0.2116 0.2192 0.2048 0.2132
MaLA-500 Llama 2 10B v2 0.2116 0.2192 0.2048 0.2132
YaYi Llama 2 7B 0.2840 0.3430 0.2835 0.2111
TowerBase Llama 2 7B 0.2794 0.3532 0.2682 0.2051
TowerInstruct Llama 2 7B 0.3010 0.3888 0.2885 0.2116
Occiglot Mistral 7B v0.1 0.2977 0.3839 0.2851 0.2103
Occiglot Mistral 7B v0.1 Instruct 0.3088 0.4029 0.2965 0.2113
BLOOM 7B 0.2365 0.2627 0.2272 0.2189
BLOOMZ 7B 0.2395 0.2694 0.2274 0.2218
mGPT 0.2024 0.2011 0.1965 0.2137
mGPT 13B 0.2176 0.2299 0.2114 0.2123
XGLM 7.5B 0.2221 0.2461 0.2105 0.2110
YaYi 7B 0.2444 0.2796 0.2329 0.2191
Llama 3 8B 0.3480 0.4243 0.3553 0.2406
Llama 3.1 8B 0.3493 0.4243 0.3589 0.2400
Gemma 7B 0.3868 0.4646 0.4047 0.2606
Gemma 2 9B 0.4415 0.5459 0.4618 0.2782
Qwen 2 7B 0.3382 0.4388 0.3264 0.2317
Qwen 1.5 7B 0.2893 0.3555 0.2814 0.2192
EMMA-500 Llama 2 7B 0.2953 0.3410 0.2982 0.2334

### v2

@@ 4.7 Summarization
The evaluation results are presented in Table 11, we observe that our model, EMMA-500, performs comparably to other Llama2-based models like TowerInstruct Llama 2 7B. Comparing with recent advances, our EMMA-500 gets slightly better performance than Llama 3 and Gemma. While EMMA-500 shows strength in certain language categories, it does not significantly outperform these models overall. This suggests that, although EMMA-500 is effective in multilingual summarization tasks, there is room for improvement to achieve more consistent performance across all languages. It is also important to note that this test set includes only 44 languages, limiting the evaluation of EMMA-500’s capabilities in low-resource languages.
Table 11: Results on XL-Sum (ROUGE-L/BERTScore %).
@@
Llama 2 7B 7.11/66.52 7.35/66.79 6.65/66.28 6.50/64.67 9.73/61.90 8.22/71.32
Llama 2 7B Chat 8.84/68.44 9.48/68.48 8.27/68.15 7.57/66.82 12.14/64.16 10.18/73.54
CodeLlama 2 7B 7.15/65.74 7.28/65.08 6.80/65.90 6.37/62.96 9.93/63.54 8.26/71.32
LLaMAX Llama 2 7B 5.29/64.59 5.07/64.83 5.19/64.71 4.94/61.43 7.19/61.85 5.94/69.14
LLaMAX Llama 2 7B Alpaca 10.11/69.24 10.42/69.70 9.41/68.68 9.45/68.14 12.65/64.86 12.15/73.80
MaLA-500 Llama 2 10B v1 5.45/63.96 5.97/64.03 5.21/63.88 4.77/61.12 7.05/63.83 5.59/68.16
MaLA-500 Llama 2 10B v2 5.44/64.28 5.86/64.45 5.22/64.07 4.86/61.35 6.86/64.33 5.60/68.82
Yayi Llama 2 7B 7.98/67.21 8.67/67.39 7.48/67.29 7.05/65.59 10.95/60.55 8.55/71.44
TowerBase Llama 2 7B 7.65/67.09 7.92/67.26 7.25/67.06 6.64/65.02 9.90/63.33 9.16/71.21
TowerInstruct Llama 2 7B 8.89/68.46 9.44/68.60 8.59/68.51 7.64/66.28 11.90/63.71 9.45/72.89
Occiglot Mistral 7B v0.1 7.33/66.20 7.53/66.40 6.82/65.87 6.80/64.31 9.71/62.74 8.77/71.11
Occiglot Mistral 7B v0.1 Instruct 8.31/66.96 9.03/66.81 7.77/67.07 7.93/63.86 10.41/63.58 8.59/72.52
BLOOM 7B 6.99/64.78 6.78/65.06 6.97/64.77 6.12/61.89 9.79/58.16 7.61/70.93
BLOOMZ 7B 11.15/69.82 13.92/68.73 8.92/69.99 10.24/69.05 12.03/66.18 14.86/74.12
mGPT 3.85/59.74 4.82/59.91 3.60/60.10 2.35/56.02 6.81/59.22 3.67/63.30
mGPT-13B 4.96/62.20 5.85/62.02 4.59/62.54 3.80/58.97 7.92/59.13 4.90/66.99
Yayi 7B 12.06/69.74 13.98/68.93 10.50/69.49 11.05/68.91 12.45/65.87 15.29/75.26
Aya 23 8B 8.68/66.79 10.03/67.66 8.24/66.26 6.80/68.73 12.63/63.26 8.47/65.68
Aya expanse 8B 10.51/68.73 11.61/68.56 9.66/68.54 10.21/71.02 12.88/65.38 10.90/68.02
Llama 3 8B 8.47/67.08 8.06/66.79 8.10/66.95 8.36/64.85 9.65/65.08 10.52/72.20
Llama 3.1 8B 8.57/66.97 8.01/66.41 8.40/66.84 8.48/65.21 9.05/65.80 10.45/71.68
Gemma 2 9B 7.38/65.45 7.24/64.00 6.87/65.81 7.47/62.95 9.22/63.57 8.89/71.46
Gemma 7B 6.70/64.52 6.86/63.48 6.24/64.91 6.63/67.21 6.73/59.86 8.30/63.35
Qwen 1.5 7B 9.58/69.13 10.61/68.78 8.58/68.89 9.61/67.83 10.79/64.92 10.74/74.40
Qwen 2 7B 10.18/69.35 11.58/69.36 9.31/68.98 9.63/68.03 10.64/64.82 11.15/74.47
EMMA-500 Llama 2 7B 8.58/67.20 8.22/65.87 8.24/67.26 8.00/69.48 11.47/65.55 10.37/67.41
@@ 4.8 Math
We evaluate the ability of LLMs to solve grade-school math problems across multiple languages on the MGSM (Multilingual Grade School Math) benchmark (113). MGSM is an extension of the GSM8K (Grade School Math 8K) dataset (26) by translating 250 of the original GSM8K problems into ten languages. We also split these ten languages into three groups, i.e., high-resource (Spanish, French, German, Russian, Chinese, and Japanese), medium-resource (Thai, Swahili, and Bengali), and low-resource (Telugu). Table 12 shows the results for 3-shot prompting with a flexible match to obtain the answers in model generation. We evaluate all the models by directly prompting the questions (denoted as direct) and the questions with answers followed by Chain-of-Thoughts prompt in the same languages as the subset being evaluated (denoted as CoT) (113). The results show that Llama 2 7B is a weak model in the MGSM math task. The base model and its variants failed in most settings. Multilingual LLMs such as BLOOM, XGLM, and YaYi are also subpar at this task. Recent LLMs like Aya, Llama 3, Qwen, and Gemma obtain reasonable performance, and Qwen series models are the best. Our model improves the Llama 2 7B model remarkably in both prompt strategies. For direct prompting, our model has an average accuracy of 0.1702, 7% higher than the Llama 2 7B chat model. For CoT-based prompting, our model increases the score of the Llama 2 7B chat model from 0.1 to 0.3 (20% higher) and slightly outperforms Llama 3 and 3.1 8B models.
Table 12: 3-shot results (ACC %) on MGSM obtained by direct and CoT prompting.
@@
Llama 2 7B 6.69 8.07 2.13 1.20 6.36 7.60 2.13 0.80
Llama 2 7B Chat 10.22 13.73 2.13 0.80 10.91 13.53 2.80 1.60
CodeLlama 2 7B 5.93 7.07 2.93 1.20 6.64 8.73 2.67 2.00
LLaMAX Llama 2 7B 3.35 4.00 2.00 0.80 3.62 4.33 2.27 2.40
LLaMAX Llama 2 7B Alpaca 5.05 5.20 4.00 1.60 6.35 8.07 4.13 0.80
MaLA-500 Llama 2 10B v1 0.91 1.33 0.27 0.00 0.73 1.27 0.00 0.00
MaLA-500 Llama 2 10B v2 0.91 1.33 0.27 0.00 0.73 1.27 0.00 0.00
YaYi Llama 2 7B 7.09 9.47 1.47 1.20 7.22 8.73 2.27 1.20
TowerBase Llama 2 7B 6.15 8.33 1.73 0.80 6.16 8.60 2.40 0.80
TowerInstruct Llama 2 7B 7.24 9.53 1.73 2.00 8.24 10.47 1.87 1.20
Occiglot Mistral 7B v0.1 13.31 16.87 4.53 1.60 14.07 18.80 3.60 1.20
Occiglot Mistral 7B v0.1 Instruct 22.76 29.80 7.47 2.80 22.16 30.40 7.87 2.80
BLOOM 7B 2.87 2.60 2.80 3.60 2.29 2.20 1.47 2.00
BLOOMZ 7B 2.55 2.67 2.40 1.20 2.15 1.67 3.07 2.00
mGPT 1.35 1.67 0.53 0.00 1.42 1.93 0.67 0.00
mGPT 13B 1.31 1.80 0.67 0.00 1.53 1.67 1.20 0.00
YaYi 7B 2.76 2.93 1.47 2.40 3.02 2.93 2.40 2.00
Aya 23 8B 22.29 30.67 3.47 0.80 24.71 35.07 5.47 2.40
Aya expanse 8B 43.02 55.47 18.93 5.20 41.45 54.67 18.53 5.20
Llama 3 8B 27.45 27.87 26.13 5.60 28.13 28.53 26.67 5.20
Llama 3.1 8B 28.36 29.00 26.13 4.40 27.31 27.27 25.47 8.40
Gemma 7B 38.22 36.60 38.27 27.20 35.78 34.67 37.07 26.80
Gemma 2 9B 32.95 28.00 35.73 30.80 44.69 36.07 52.00 47.20
Qwen 2 7B 48.95 54.40 38.80 14.80 51.47 58.93 39.07 14.40
Qwen 1.5 7B 31.56 40.00 16.00 4.00 30.36 40.60 14.80 2.40
EMMA-500 Llama 2 7B 17.02 19.20 11.87 2.40 18.09 20.00 13.20 2.80
@@ 4.9 Machine Reading Comprehension
BELEBELE (16) is a machine reading comprehension dataset covering 122 languages including high- and low-resource languages. Each question offers four multiple-choice answers based on a short passage from the FLORES-200 dataset. This benchmark is very challenging, even the English version of it presents remarkable challenges for advanced models. Table 13 shows the zero-shot results in different resource groups. Our continual pre-training improves the Llama 2 7B base model. But Llama 2 7 B-based models mostly fail in this task and get quasi-random results. Recent advanced models like Aya expanse, Llama 3.1 and Qwen 2 get reasonable results.
Table 13: 0-shot results (ACC %) on BELEBELE.
@@
Llama 2 7B 26.27 26.76 26.35 26.07 26.41 26.27
Llama 2 7B Chat 29.05 31.84 29.97 28.95 29.47 29.09
CodeLlama 2 7B 27.38 27.37 27.33 27.30 27.30 27.38
LLaMAX Llama 2 7B 23.09 23.23 23.15 23.07 23.10 23.08
LLaMAX Llama 2 7B Alpaca 24.48 25.46 24.82 24.41 24.60 24.49
MaLA-500 Llama 2 10B v1 22.96 23.02 22.98 22.97 22.98 22.97
MaLA-500 Llama 2 10B v2 22.96 23.02 22.98 22.97 22.98 22.97
YaYi Llama 2 7B 28.32 29.64 28.67 28.11 28.37 28.26
TowerBase Llama 2 7B 26.36 27.43 26.85 26.29 26.48 26.34
TowerInstruct Llama 2 7B 27.93 29.88 28.51 27.57 28.19 27.92
Occiglot Mistral 7B v0.1 30.16 32.25 30.94 30.02 30.40 30.15
Occiglot Mistral 7B v0.1 Instruct 32.05 34.14 32.62 31.74 32.40 32.08
BLOOM 7B 24.11 24.25 24.52 24.12 24.11 24.08
BLOOMZ 7B 39.32 45.43 43.67 41.51 40.08 39.51
mGPT 23.96 24.01 23.81 24.00 23.94 23.98
YaYi 7B 37.97 44.37 42.71 40.49 38.72 38.09
Aya 23 8B 40.08 43.85 41.71 39.22 40.93 39.81
Aya expanse 8B 46.98 52.22 48.99 46.57 48.36 46.93
Llama 3 8B 40.73 46.07 42.92 41.03 41.87 40.88
Llama 3.1 8B 45.19 52.50 48.01 45.65 46.74 45.34
Gemma 7B 43.37 52.63 47.83 44.82 45.43 43.94
Gemma 2 9B 54.49 64.10 58.90 55.83 56.85 55.05
Qwen 2 7B 49.31 57.62 52.20 50.04 51.16 49.48
Qwen 1.5 7B 41.83 48.86 44.79 42.18 43.00 41.78
EMMA-500 Llama 2 7B 26.75 28.32 28.18 27.58 27.14 26.94
@@
Table 14: 5-shot results (ACC %) on ARC multilingual.
@@
Llama 2 7B 27.5608 33.1244 27.3065 21.0196
Llama 2 7B Chat 28.0155 33.6893 27.7891 21.2910
CodeLlama 2 7B 25.2283 28.8639 24.6368 21.6450
LLaMAX Llama 2 7B 26.0905 30.0022 25.9174 21.4821
LLaMAX Llama 2 7B Alpaca 31.0602 36.8898 31.8518 22.4866
MaLA-500 Llama 2 10B v1 21.1612 21.9174 20.4836 21.3172
MaLA-500 Llama 2 10B v2 21.1612 21.9174 20.4836 21.3172
YaYi Llama 2 7B 28.3975 34.2968 28.3460 21.1068
TowerBase Llama 2 7B 27.9355 35.3215 26.8248 20.5081
TowerInstruct Llama 2 7B 30.1019 38.8803 28.8544 21.1562
Occiglot Mistral 7B v0.1 29.7667 38.3930 28.5078 21.0296
Occiglot Mistral 7B v0.1 Instruct 30.8844 40.2927 29.6522 21.1264
BLOOM 7B 23.6501 26.2685 22.7178 21.8920
BLOOMZ 7B 23.9497 26.9360 22.7439 22.1762
mGPT 20.2422 20.1129 19.6471 21.3709
mGPT 13B 21.7584 22.9872 21.1366 21.2329
YaYi 7B 24.4413 27.9617 23.2904 21.9109
Aya 23 8B 31.0838 40.0507 30.0173 21.6080
Aya expanse 8B 36.5617 47.8696 36.1507 23.0948
Llama 3 8B 34.7957 42.4323 35.5256 24.0638
Llama 3.1 8B 34.9336 42.4323 35.8939 23.9995
Gemma 7B 38.6827 46.4619 40.4660 26.0609
Gemma 2 9B 44.1534 54.5890 46.1756 27.8228
Qwen 2 7B 33.8204 43.8754 32.6423 23.1660
Qwen 1.5 7B 28.9263 35.5527 28.1392 21.9223
EMMA-500 Llama 2 7B 29.5282 34.1004 29.8165 23.3444